Complete Publication List of Philippe Flajolet

[1] Philippe Flajolet and Jean-Marc Steyaert, Decision problems for multihead finite automata, Mathematical foundations of computer science (Proc. Sympos. and Summer School, Štrbské Pleso, 1973), Math. Inst., Slovak Acad. Sci., Bratislava, 1973, pp. 225–230. MR0408315 (53 #12080)

[2] P. Flajolet and J. M. Steyaert, Complexité des problèmes de décision relatifs aux algorithmes de tri, Automata, languages and programming (Proc. Sympos., Rocquencourt, 1972), North-Holland, Amsterdam, 1973, pp. 537–548 (French). MR0416107 (54 #4183)

[3] Philippe Flajolet and Jean-Marc Steyaert, Une généralisation de la notion d’ensemble immune, Rev. Française Automat. Informat. Recherche Opérationnelle Sér. Rouge 8 (1974), no. R-1, 37–48 (French). MR0349364 (50 #1858)

[4] P. Flajolet and J. M. Steyaert, On sets having only hard subsets, Automata, languages and programming (Second Colloq., Univ. Saarbrücken, Saarbrücken, 1974), Springer, Berlin, 1974, pp. 446–457. Lecture Notes in Comput. Sci., Vol. 14. MR0424541 (54 #12501)

[5] P. Flajolet and J. M Steyaert, Hiérarchies de complexité et réductions entre problèmes, Journées algorithmiques (École Norm. Sup., Paris, 1975), Soc. Math. France, Paris, 1976, pp. 53–72. Astérisque, No. 38-39 (French). MR0457164 (56 #15380)

[6] P. Flajolet, J. C. Raoult, and J. Vuillemin, On the average number of registers required for evaluating arithmetic expressions, 18th Annual Symposium on Foundations of Computer Science (Providence, R.I., 1977), IEEE Comput. Sci., Long Beach, Calif., 1977, pp. 196–205. MR0495192 (58 #13923)

[7] Philippe Flajolet, Analyse en moyenne de la détection des arbres partiels, Les arbres en algèbre et en programmation (Troisième Colloq. de Lille, Lille, 1978), Univ. Sci. Tech. Lille, Villeneuve d’Ascq, 1978, pp. 134–138 (French). MR505846 (80g:68087)

[8] P. Flajolet, J.-C. Raoult, and J. Vuillemin, The number of registers required for evaluating arithmetic expressions, Theoret. Comput. Sci. 9 (1979), no. 1, 99–125. MR535127 (80e:68101)

[9] Philippe Flajolet, Analyse d’algorithmes de manipulation de fichiers, Les arbres en algèbre et en programmation (4ème Colloq., Lille, 1979), Univ. Lille I, Lille, 1979, pp. 296–318 (French). MR554354 (81c:68024)

[10] P. Flajolet, J. Françon, and J. Vuillemin, Computing integrated costs of sequences of operations with application to dictionaries, Computing (Atlanta, Ga., 1979), ACM, New York, 1979, pp. 49–61. MR564619 (81g:68027)

[11] P. Flajolet and Lyle Ramshaw, A note on Gray code and odd-even merge, SIAM J. Comput. 9 (1980), no. 1, 142–158. MR557835 (81h:68025)

[12] Philippe Flajolet, Deux problèmes d’analyse d’algorithmes, Séminaire Delange-Pisot-Poitou, 20e année: 1978/1979. Théorie des nombres, Fasc. 1 (French), Secrétariat Math., Paris, 1980, pp. Exp. No. 14, 10 (French). MR582426 (81i:68059)

[13] Philippe Flajolet and Jean-Marc Steyaert, On the analysis of tree-matching algorithms, Automata, languages and programming (Proc. Seventh Internat. Colloq., Noordwijkerhout, 1980), Lecture Notes in Comput. Sci., vol. 85, Springer, Berlin, 1980, pp. 208–219. MR589005 (82d:68023)

[14] P. Flajolet, Combinatorial aspects of continued fractions, Discrete Math. 32 (1980), no. 2, 125–161. MR592851 (82f:05002a)

[15] Philippe Flajolet, Combinatorial aspects of continued fractions, Ann. Discrete Math. 9 (1980), 217–222. Combinatorics 79 (Proc. Colloq., Univ. Montréal, Montreal, Que., 1979), Part II. MR597375 (82f:05002b)

[16] Philippe Flajolet and Jean-Marc Steyaert, On the analysis of tree-matching algorithms, Trees in algebra and programming (Proc. 5th Lille Colloq., Lille, 1980), Univ. Lille I, Lille, 1980, pp. 22–40. MR620177 (82g:68030)

[17] P. Flajolet, J. Françon, and J. Vuillemin, Sequence of operations analysis for dynamic data structures, J. Algorithms 1 (1980), no. 2, 111–141. MR604861 (82m:68039)

[18] Philippe Flajolet and Jean-Marc Steyaert, A branching process arising in dynamic hashing, trie searching and polynomial factorization, Automata, languages and programming (Aarhus, 1982), Lecture Notes in Comput. Sci., vol. 140, Springer, Berlin, 1982, pp. 239– 251. MR675461 (83m:68108)

[19] Philippe Flajolet and Andrew Odlyzko, The average height of binary trees and other simple trees, J. Comput. System Sci. 25 (1982), no. 2, 171–213. MR680517 (84a:68056)

[20] Philippe Flajolet, On congruences and continued fractions for some classical combinatorial quantities, Discrete Math. 41 (1982), no. 2, 145–153. MR676874 (84f:05005)

[21] , On the performance evaluation of extendible hashing and trie searching, Acta Inform. 20 (1983), no. 4, 345–369. MR732311 (85a:68022)

[22] , Methods in the analysis of algorithms: evaluations of a recursive partitioning process, Foundations of computation theory (Borgholm, 1983), Lecture Notes in Comput. Sci., vol. 158, Springer, Berlin, 1983, pp. 141–158. MR734716 (85g:68013)

[23] Philippe Flajolet and Nasser Saheb, Digital search trees and the generation of an exponentially distributed variate, CAAP ’83 (L’Aquila, 1983), Lecture Notes in Comput. Sci., vol. 159, Springer, Berlin, 1983, pp. 221–235. MR744212 (85e:68029)

[24] Jean-Marc Steyaert and Philippe Flajolet, Patterns and pattern-matching in trees: an analysis, Inform. and Control 58 (1983), no. 1-3, 19–58. MR750401 (86h:68070)

[25] P. Flajolet and A. M. Odlyzko, Limit distributions for coefficients of iterates of polynomials with applications to combinatorial enumerations, Math. Proc. Cambridge Philos. Soc. 96 (1984), no. 2, 237–253. MR757658 (86j:68105)

[26] Philippe Flajolet and Jean-Marc Steyaert, A complexity calculus for recursive tree algorithms, (Avignon, 1983), Publ. Dép. Math. Nouvelle Sér. B, vol. 84, Univ. Claude-Bernard, Lyon, 1984, pp. 39–88 (English, with French summary). MR802632 (86k:68040)

[27] Philippe Flajolet, Approximate counting: a detailed analysis, BIT 25 (1985), no. 1, 113–134. MR785808 (86j:68053)

[28] Peter Mathys and Philippe Flajolet, Q-ary collision resolution algorithms in random-access systems with free or blocked channel access, IEEE Trans. Inform. Theory 31 (1985), no. 2, 217–243. MR793093 (86k:94026)

[29] Guy Fayolle, Philippe Flajolet, Micha Hofri, and Philippe Jacquet, Analysis of a stack algorithm for random multiple-access communication, IEEE Trans. Inform. Theory 31 (1985), no. 2, 244–254. MR793094 (86g:94015)

[30] Philippe Flajolet, Thomas Ottmann, and Derick Wood, Search trees and bubble memories, RAIRO Inform. Théor. 19 (1985), no. 2, 137–164 (English, with French summary). MR803995 (86m:68014)

[31] Philippe Flajolet, Mireille Régnier, and Dominique Sotteau, Algebraic methods for trie statistics, Analysis and design of algorithms for combinatorial problems (Udine, 1982), North-Holland Math. Stud., vol. 109, North-Holland, Amsterdam, 1985, pp. 145–188. MR808001 (86m:68013)

[32] Philippe Flajolet, Mireille Régnier, and Robert Sedgewick, Some uses of the Mellin integral transform in the analysis of algorithms, Combinatorial algorithms on words (Maratea, 1984), NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., vol. 12, Springer, Berlin, 1985, pp. 241– 254. MR815343 (86m:44001)

[33] Philippe Flajolet, Ambiguity and transcendence, Automata, languages and programming (Nafplion, 1985), Lecture Notes in Comput. Sci., vol. 194, Springer, Berlin, 1985, pp. 179– 188. MR819253 (87c:68041)

[34] , Elements of a general theory of combinatorial structures, Fundamentals of computation theory (Cottbus, 1985), Lecture Notes in Comput. Sci., vol. 199, Springer, Berlin, 1985, pp. 112–127. MR821230 (87e:68040)

[35] Philippe Flajolet and G. Nigel Martin, Probabilistic counting algorithms for data base applications, J. Comput. System Sci. 31 (1985), no. 2, 182–209. Special issue: Twenty-fourth annual symposium on the foundations of computer science (Tucson, Ariz., 1983). MR828521 (87h:68023)

[36] Philippe Flajolet and Claude Puech, Partial match retrieval of multidimensional data, J. Assoc. Comput. Mach. 33 (1986), no. 2, 371–407. MR835110 (87j:68024)

[37] Philippe Flajolet, Claude Puech, and Jean Vuillemin, The analysis of simple list structures, Inform. Sci. 38 (1986), no. 2, 121–146. MR835816 (87f:68010)

[38] Guy Fayolle, Philippe Flajolet, and Micha Hofri, On a functional equation arising in the analysis of a protocol for a multi-access broadcast channel, Adv. in Appl. Probab. 18 (1986), no. 2, 441–472. MR840103 (87g:68008)

[39] P. Flajolet and H. Prodinger, Register allocation for unary-binary trees, SIAM J. Comput. 15 (1986), no. 3, 629–640. MR850413 (87j:68052)

[40] Philippe Flajolet and Robert Sedgewick, Digital search trees revisited, SIAM J. Comput. 15 (1986), no. 3, 748–767. MR850421 (87m:68014)

[41] Philippe Flajolet and Nasser Saheb, The complexity of generating an exponentially distributed variate, J. Algorithms 7 (1986), no. 4, 463–488. MR866787 (88b:65012)

[42] Philippe Flajolet, The evolution of two stacks in bounded space and random walks in a triangle, Mathematical foundations of computer science, 1986 (Bratislava, 1986), Lecture Notes in Comput. Sci., vol. 233, Springer, Berlin, 1986, pp. 325–340. MR874610

[43] Philippe Flajolet and Jean-Marc Steyaert, A complexity calculus for recursive tree algorithms, Math. Systems Theory 19 (1987), no. 4, 301–331. MR888494 (88h:68033)

[44] Albert G. Greenberg, Philippe Flajolet, and Richard E. Ladner, Estimating the multiplicities of conflicts to speed their resolution in multiple access channels, J. Assoc. Comput. Mach. 34 (1987), no. 2, 289–325. MR892225 (88f:68010)

[45] Philippe Flajolet and Helmut Prodinger, Level number sequences for trees, Discrete Math. 65 (1987), no. 2, 149–156. MR893076 (88e:05030)

[46] Jean-Michel Autebert, Philippe Flajolet, and Joaquim Gabarró, Prefixes of infinite words and ambiguous context-free languages, Inform. Process. Lett. 25 (1987), no. 4, 211–216. MR896136 (88m:68027)

[47] Philippe Flajolet, Analytic models and ambiguity of context-free languages, Theoret. Comput. Sci. 49 (1987), no. 2-3, 283–309. Twelfth international colloquium on automata, languages and programming (Nafplion, 1985). MR909335 (89e:68067)

[48] , Mathematical methods in the analysis of algorithms and data structures, Trends in theoretical computer science (Udine, 1984), Principles Comput. Sci. Ser., vol. 12, Computer Sci. Press, Rockville, MD, 1988, pp. 225–304. MR918993

[49] P. Flajolet, P. Kirschenhofer, and R. F. Tichy, Deviations from uniformity in random strings, Probab. Theory Related Fields 80 (1988), no. 1, 139–150. MR970475 (90a:11087)

[50] Philippe Flajolet, Danièle Gardy, and Loÿs Thimonier, Random allocations and probabilistic languages, Automata, languages and programming (Tampere, 1988), Lecture Notes in Comput. Sci., vol. 317, Springer, Berlin, 1988, pp. 239–253. MR1023640 (91a:68127)

[51] Philippe Flajolet, Random tree models in the analysis of algorithms, Performance ’87 (Brussels, 1987), North-Holland, Amsterdam, 1988, pp. 171–187. MR1024826 (90k:68067)

[52] Ph. Flajolet, P. Kirschenhofer, and R. F. Tichy, Discrepancy of sequences in discrete spaces, Irregularities of partitions (Fertőd, 1986), Algorithms Combin. Study Res. Texts, vol. 8, Springer, Berlin, 1989, pp. 61–70. MR999931 (90g:11105)

[53] Philippe Flajolet, Donald E. Knuth, and Boris Pittel, The first cycles in an evolving graph, Discrete Math. 75 (1989), no. 1-3, 167–215. Graph theory and combinatorics (Cambridge, 1988). MR1001395 (90d:05184)

[54] Philippe Flajolet, Bruno Salvy, and Paul Zimmermann, Lambda-upsilon-omega: an assistant algorithms analyzer, Applied algebra, algebraic algorithms and error-correcting codes (Rome, 1988), Lecture Notes in Comput. Sci., vol. 357, Springer, Berlin, 1989, pp. 201–212. MR1008504

[55] Philippe Flajolet and Jean Françon, Elliptic functions, continued fractions and doubled permutations, European J. Combin. 10 (1989), no. 3, 235–241. MR1029169 (90k:05010)

[56] Walter Cunto, Gustavo Lau, and Philippe Flajolet, Analysis of kdt-trees: kd-trees improved by local reorganisations, Algorithms and data structures (Ottawa, ON, 1989), Lecture Notes in Comput. Sci., vol. 382, Springer, Berlin, 1989, pp. 24–38. MR1035763

[57] Danièle Gardy, Philippe Flajolet, and Claude Puech, On the performance of orthogonal range queries in multiattribute and doubly chained trees, Algorithms and data structures (Ottawa, ON, 1989), Lecture Notes in Comput. Sci., vol. 382, Springer, Berlin, 1989, pp. 218– 229. MR1035772

[58] Philippe Flajolet and Andrew Odlyzko, Singularity analysis of generating functions, SIAM J. Discrete Math. 3 (1990), no. 2, 216–240. MR1039294 (90m:05012)

[59] Philippe Flajolet and Michèle Soria, Gaussian limiting distributions for the number of components in combinatorial structures, J. Combin. Theory Ser. A 53 (1990), no. 2, 165–182. MR1041444 (91c:05012)

[60] P. Flajolet, On adaptive sampling, Computing 43 (1990), no. 4, 391–400 (English, with German summary). MR1045070 [61] J.-P. Allouche, P. Flajolet, and M. Mendès France, Algebraically independent formal power

series: a language theory interpretation, Analytic number theory (Tokyo, 1988), Lecture Notes in Math., vol. 1434, Springer, Berlin, 1990, pp. 11–18. MR1071741 (91j:11013)

[62] Philippe Flajolet and René Schott, Nonoverlapping partitions, continued fractions, Bessel functions and a divergent series, European J. Combin. 11 (1990), no. 5, 421–432. MR1075531 (91m:05015)

[63] Philippe Flajolet, Paolo Sipala, and Jean-Marc Steyaert, Analytic variations on the common subexpression problem, Automata, languages and programming (Coventry, 1990), Lecture Notes in Comput. Sci., vol. 443, Springer, New York, 1990, pp. 220–234. MR1076824

[64] Philippe Flajolet and Andrew M. Odlyzko, Random mapping statistics, Advances in cryptology—EUROCRYPT ’89 (Houthalen, 1989), Lecture Notes in Comput. Sci., vol. 434, Springer, Berlin, 1990, pp. 329–354. MR1083961

[65] Jeffrey Scott Vitter and Philippe Flajolet, Average-case analysis of algorithms and data structures, Handbook of theoretical computer science, Vol. A, Elsevier, Amsterdam, 1990, pp. 431–524. MR1127175

[66] Brigitte Vallée and Philippe Flajolet, The lattice reduction algorithm of Gauss: an average case analysis, (St. Louis, MO, 1990), IEEE Comput. Soc. Press, Los Alamitos, CA, 1990, pp. 830–839. MR1150732 (92m:11153)

[67] P. Flajolet and M. Soria, The cycle construction, SIAM J. Discrete Math. 4 (1991), no. 1, 58–60. MR1090289 (92j:05005)

[68] Philippe Flajolet, Gaston Gonnet, Claude Puech, and J. M. Robson, The analysis of multidimensional searching in quad-trees, Algorithms (San Francisco, CA, 1991), ACM, New York, 1991, pp. 100–109. MR1095824

[69] Philippe Flajolet, Bruno Salvy, and Paul Zimmermann, Automatic average-case analysis of algorithms. no. 1 (Part A), Theoret. Comput. Sci. 79 (1991), no. 1, (Part A), 37–109. Algebraic and computing treatment of noncommutative power series (Lille, 1988). MR1102951 (92k:68049)

[70] Philippe Flajolet and Bruce Richmond, Generalized digital trees and their differencedifferential equations, Random Structures Algorithms 3 (1992), no. 3, 305–320. MR1164843 (93f:05086)

[71] Mamoru Hoshi and Philippe Flajolet, Page usage in a quadtree index, BIT 32 (1992), no. 3, 384–402. MR1179228 (94b:68040)

[72] Philippe Flajolet, Danièle Gardy, and Loÿs Thimonier, Birthday paradox, coupon collectors, caching algorithms and self-organizing search, Discrete Appl. Math. 39 (1992), no. 3, 207– 229. MR1189469 (93i:68107)

[73] Philippe Flajolet, Analytic analysis of algorithms, Automata, languages and programming (Vienna, 1992), Lecture Notes in Comput. Sci., vol. 623, Springer, Berlin, 1992, pp. 186–210. MR1250638

[74] François Bergeron, Philippe Flajolet, and Bruno Salvy, Varieties of increasing trees, CAAP ’92 (Rennes, 1992), Lecture Notes in Comput. Sci., vol. 581, Springer, Berlin, 1992, pp. 24– 48. MR1251994 (94j:68233)

[75] Philippe Flajolet and Michèle Soria, General combinatorial schemas: Gaussian limit distributions and exponential tails, Discrete Math. 114 (1993), no. 1-3, 159–180. Combinatorics and algorithms (Jerusalem, 1988). MR1217750 (94e:05021)

[76] Philippe Flajolet, Xavier Gourdon, and Bruno Salvy, Sur une famille de polynômes issus de l’analyse numérique, Gaz. Math. 55 (1993), 67–78 (French). MR1244460 (95c:65036)

[77] Philippe Flajolet, Gaston Gonnet, Claude Puech, and J. M. Robson, Analytic variations on quadtrees, Algorithmica 10 (1993), no. 6, 473–500. MR1244619 (94i:68052)

[78] Philippe Flajolet, Zhicheng Gao, Andrew Odlyzko, and Bruce Richmond, The distribution of heights of binary trees and other simple trees, Combin. Probab. Comput. 2 (1993), no. 2, 145–156. MR1249127 (94k:05061)

[79] Philippe Flajolet and Mordecai Golin, Exact asymptotics of divide-and-conquer recurrences, Automata, languages and programming (Lund, 1993), Lecture Notes in Comput. Sci., vol. 700, Springer, Berlin, 1993, pp. 137–149. MR1252408

[80] Philippe Flajolet, Paul Zimmermann, and Bernard Van Cutsem, A calculus of random generation, Algorithms—ESA ’93 (Bad Honnef, 1993), Lecture Notes in Comput. Sci., vol. 726, Springer, Berlin, 1993, pp. 169–180. MR1284724 [81] Philippe Flajolet, Peter Grabner, Peter Kirschenhofer, Helmut Prodinger, and Robert F.

Tichy, Mellin transforms and asymptotics: digital sums, Theoret. Comput. Sci. 123 (1994), no. 2, 291–314. MR1256203 (94m:11090)

[82] P. Flajolet and T. Lafforgue, Search costs in quadtrees and singularity perturbation asymptotics, Discrete Comput. Geom. 12 (1994), no. 2, 151–175. MR1283884 (95d:68020)

[83] Philippe Flajolet, Paul Zimmerman, and Bernard Van Cutsem, A calculus for the random generation of labelled combinatorial structures, Theoret. Comput. Sci. 132 (1994), no. 1-2, 1–35. MR1290534 (96f:05172)

[84] Philippe Flajolet and Mordecai Golin, Mellin transforms and asymptotics. The mergesort recurrence, Acta Inform. 31 (1994), no. 7, 673–696. MR1300060 (95h:68035)

[85] Hervé Daudé, Philippe Flajolet, and Brigitte Vallée, An analysis of the Gaussian algorithm for lattice reduction, Algorithmic number theory (Ithaca, NY, 1994), Lecture Notes in Comput. Sci., vol. 877, Springer, Berlin, 1994, pp. 144–158. MR1322719 (96a:11075)

[86] Philippe Flajolet, Xavier Gourdon, and Philippe Dumas, Mellin transforms and asymptotics: harmonic sums, Theoret. Comput. Sci. 144 (1995), no. 1-2, 3–58. Special volume on mathematical analysis of algorithms. MR1337752 (96h:68093)

[87] Philippe Flajolet and Robert Sedgewick, Mellin transforms and asymptotics: finite differences and Rice’s integrals, Theoret. Comput. Sci. 144 (1995), no. 1-2, 101–124. Special volume on mathematical analysis of algorithms. MR1337755 (96i:39003)

[88] Philippe Flajolet, Peter J. Grabner, Peter Kirschenhofer, and Helmut Prodinger, On Ramanujan’s Q-function, J. Comput. Appl. Math. 58 (1995), no. 1, 103–116. MR1344359 (96f:65010)

[89] Philippe Flajolet, Gilbert Labelle, Louise Laforest, and Bruno Salvy, Hypergeometrics and the cost structure of quadtrees, Random Structures Algorithms 7 (1995), no. 2, 117–144. MR1369059 (96m:68034)

[90] Philippe Flajolet and Bruno Salvy, Computer algebra libraries for combinatorial structures, J. Symbolic Comput. 20 (1995), no. 5-6, 653–671. Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993). MR1395419 (97j:68096)

[91] Philippe Dumas and Philippe Flajolet, Asymptotique des récurrences mahlériennes: le cas cyclotomique, J. Théor. Nombres Bordeaux 8 (1996), no. 1, 1–30 (French, with English and French summaries). MR1399944 (97f:39029)

[92] Philippe Flajolet, Xavier Gourdon, and Daniel Panario, Random polynomials and polynomial factorization, Automata, languages and programming (Paderborn, 1996), Lecture Notes in Comput. Sci., vol. 1099, Springer, Berlin, 1996, pp. 232–243. MR1464452 (98e:68123)

[93] Hervé Daudé, Philippe Flajolet, and Brigitte Vallée, An average-case analysis of the Gaussian algorithm for lattice reduction, Combin. Probab. Comput. 6 (1997), no. 4, 397–433. MR1483426 (99a:65196)

[94] Zhen Yu Wang and Ph. Flajolet, Some enumeration characteristics of recursive trees, J. Systems Sci. Math. Sci. 17 (1997), no. 3, 198–203 (Chinese, with English and Chinese summaries). MR1490181 (98j:05053)

[95] Philippe Flajolet, Xavier Gourdon, and Conrado Martínez, Patterns in random binary search trees, Random Structures Algorithms 11 (1997), no. 3, 223–244. MR1609509 (98m:68042)

[96] Philippe Flajolet and Brigitte Vallée, Continued fraction algorithms, functional operators, and structure constants, Theoret. Comput. Sci. 194 (1998), no. 1-2, 1–34. MR1491644 (98j:11061)

[97] Philippe Flajolet and Bruno Salvy, Euler sums and contour integral representations, Experiment. Math. 7 (1998), no. 1, 15–35. MR1618286 (99c:11110)

[98] E. G. Coffman Jr., Philippe Flajolet, Leopold Flatto, and Micha Hofri, The maximum of a random walk and its application to rectangle packing, Probab. Engrg. Inform. Sci. 12 (1998), no. 3, 373–386. MR1631315 (99f:60127)

[99] Julien Clément, Philippe Flajolet, and Brigitte Vallée, The analysis of hybrid trie structures, Algorithms (San Francisco, CA, 1998), ACM, New York, 1998, pp. 531–539. MR1642966

[100] P. Flajolet, P. Poblete, and A. Viola, On the analysis of linear probing hashing, Algorithmica 22 (1998), no. 4, 490–515. Average-case analysis of algorithms. MR1701625 (2000h:68056)

[101] Daniel Panario, Xavier Gourdon, and Philippe Flajolet, An analytic approach to smooth polynomials over finite fields, Algorithmic number theory (Portland, OR, 1998), Lecture Notes in Comput. Sci., vol. 1423, Springer, Berlin, 1998, pp. 226–236. MR1726074 (2001e:11119)

[102] Philippe Flajolet, Singularity analysis and asymptotics of Bernoulli sums, Theoret. Comput. Sci. 215 (1999), no. 1-2, 371–381. MR1678788 (2000a:05015)

[103] Philippe Flajolet and Helmut Prodinger, On Stirling numbers for complex arguments and Hankel contours, SIAM J. Discrete Math. 12 (1999), no. 2, 155–159 (electronic). MR1685895 (2000d:05005)

[104] L. Devroye, P. Flajolet, F. Hurtado, M. Noy, and W. Steiger, Properties of random triangulations and trees, Discrete Comput. Geom. 22 (1999), no. 1, 105–117. MR1692686 (2000h:52005)

[105] Philippe Flajolet and Marc Noy, Analytic combinatorics of non-crossing configurations, Discrete Math. 204 (1999), no. 1-3, 203–229. MR1691870 (2000c:05012)

[106] Pierre Nicodème, Bruno Salvy, and Philippe Flajolet, Motif statistics, Algorithms—ESA ’99 (Prague), Lecture Notes in Comput. Sci., vol. 1643, Springer, Berlin, 1999, pp. 194–211. MR1729125

[107] Hosam Mahmoud, Philippe Flajolet, Philippe Jacquet, and Mireille Régnier, Analytic variations on bucket selection and sorting, Acta Inform. 36 (2000), no. 9-10, 735–760. MR1757914 (2001f:68026)

[108] Philippe Flajolet and Brigitte Vallée, Continued fractions, comparison algorithms, and fine structure constants, Constructive, experimental, and nonlinear analysis (Limoges, 1999), CMS Conf. Proc., vol. 27, Amer. Math. Soc., Providence, RI, 2000, pp. 53–82. MR1777617 (2001h:11161)

[109] Philippe Flajolet and Fabrice Guillemin, The formal theory of birth-and-death processes, lattice path combinatorics and continued fractions, Adv. in Appl. Probab. 32 (2000), no. 3, 750–778. MR1788094 (2002f:60165)

[110] Cyril Banderier, Philippe Flajolet, Gilles Schaeffer, and Michèle Soria, Planar maps and Airy phenomena, Automata, languages and programming (Geneva, 2000), Lecture Notes in Comput. Sci., vol. 1853, Springer, Berlin, 2000, pp. 388–402. MR1795907 (2001h:68113)

[111] Philippe Flajolet and Marc Noy, Analytic combinatorics of chord diagrams, Formal power series and algebraic combinatorics (Moscow, 2000), Springer, Berlin, 2000, pp. 191–201. MR1798213 (2002i:05004)

[112] P. Flajolet, X. Gourdon, and D. Panario, The complete analysis of a polynomial factorization algorithm over finite fields, J. Algorithms 40 (2001), no. 1, 37–81. MR1841252 (2002f:68193)

[113] P. Flajolet and G. Louchard, Analytic variations on the Airy distribution, Algorithmica 31 (2001), no. 3, 361–377. Mathematical analysis of algorithms. MR1855255 (2002j:68115)

[114] Philippe Flajolet, D · E · K = (100)_{8}, Random Structures Algorithms 19 (2001), no. 3-4, 150–162. Analysis of algorithms (Krynica Morska, 2000). MR1871553 (2003d:01037)

[115] Cyril Banderier, Philippe Flajolet, Gilles Schaeffer, and Michèle Soria, Random maps, coalescing saddles, singularity analysis, and Airy phenomena, Random Structures Algorithms 19 (2001), no. 3-4, 194–246. Analysis of algorithms (Krynica Morska, 2000). MR1871555 (2002k:05012)

[116] J. Clément, P. Flajolet, and B. Vallée, Dynamical sources in information theory: a general analysis of trie structures, Algorithmica 29 (2001), no. 1-2, 307–369. Average-case analysis of algorithms (Princeton, NJ, 1998). MR1887308 (2003b:68033)

[117] Philippe Flajolet, Yves Guivarc’h, Wojciech Szpankowski, and Brigitte Vallée, Hidden pattern statistics, Automata, languages and programming, Lecture Notes in Comput. Sci., vol. 2076, Springer, Berlin, 2001, pp. 152–165. MR2065859

[118] Cyril Banderier, Mireille Bousquet-Mélou, Alain Denise, Philippe Flajolet, Danièle Gardy, and Dominique Gouyou-Beauchamps, Generating functions for generating trees, Discrete Math. 246 (2002), no. 1-3, 29–55. Formal power series and algebraic combinatorics (Barcelona, 1999). MR1884885 (2003c:05008)

[119] Cyril Banderier and Philippe Flajolet, Basic analytic combinatorics of directed lattice paths, Theoret. Comput. Sci. 281 (2002), no. 1-2, 37–80. Selected papers in honour of Maurice Nivat. MR1909568 (2003g:05006)

[120] P. Flajolet, K. Hatzis, S. Nikoletseas, and P. Spirakis, On the robustness of interconnections in random graphs: a symbolic approach, Theoret. Comput. Sci. 287 (2002), no. 2, 515–534. Algorithms (Prague, 1999). MR1930233 (2003h:05168)

[121] Pierre Nicodème, Bruno Salvy, and Philippe Flajolet, Motif statistics, Theoret. Comput. Sci. 287 (2002), no. 2, 593–617. Algorithms (Prague, 1999). MR1930238 (2003i:68104)

[122] Philippe Flajolet and Wojciech Szpankowski, Analytic variations on redundancy rates of renewal processes, IEEE Trans. Inform. Theory 48 (2002), no. 11, 2911–2921. MR1945582 (2004m:94012)

[123] Philippe Flajolet, Singular combinatorics, (Beijing, 2002), Higher Ed. Press, Beijing, 2002, pp. 561–571. MR1957561 (2003k:05009)

[124] Philippe Duchon, Philippe Flajolet, Guy Louchard, and Gilles Schaeffer, Random sampling from Boltzmann principles, Automata, languages and programming, Lecture Notes in Comput. Sci., vol. 2380, Springer, Berlin, 2002, pp. 501–513. MR2062483

[125] Philippe Chassaing and Philippe Flajolet, Hachage, arbres, chemins & graphes, Gaz. Math. 95 (2003), 29–49 (French). MR1959899 (2003k:68132)

[126] Philippe Flajolet, Bruno Salvy, and Gilles Schaeffer, Airy phenomena and analytic combinatorics of connected graphs, Electron. J. Combin. 11 (2004), no. 1, Research Paper 34, 30 pp. (electronic). MR2056086 (2005a:05014)

[127] B. Chauvin, P. Flajolet, D. Gardy, and B. Gittenberger, And/or trees revisited, Combin. Probab. Comput. 13 (2004), no. 4-5, 475–497. MR2095972 (2005g:68049)

[128] Philippe Duchon, Philippe Flajolet, Guy Louchard, and Gilles Schaeffer, Boltzmann samplers for the random generation of combinatorial structures, Combin. Probab. Comput. 13 (2004), no. 4-5, 577–625. MR2095975 (2005k:05030)

[129] Philippe Flajolet, Stefan Gerhold, and Bruno Salvy, On the non-holonomic character of logarithms, powers, and the nth prime function, Electron. J. Combin. 11 (2004/06), no. 2, Article 2, 16. MR2195433 (2006h:05013)

[130] James Allen Fill, Philippe Flajolet, and Nevin Kapur, Singularity analysis, Hadamard products, and tree recurrences, J. Comput. Appl. Math. 174 (2005), no. 2, 271–313. MR2106441 (2005h:05012)

[131] Philippe Flajolet, Joaquim Gabarró, and Helmut Pekari, Analytic urns, Ann. Probab. 33 (2005), no. 3, 1200–1233. MR2135318 (2006m:60014)

[132] Eric van Fossen Conrad and Philippe Flajolet, The Fermat cubic, elliptic functions, continued fractions, and a combinatorial excursion, Sém. Lothar. Combin. 54 (2005/07), Art. B54g, 44. MR2223029 (2007e:33020)

[133] Alin Bostan, Philippe Flajolet, Bruno Salvy, and Éric Schost, Fast computation of special resultants, J. Symbolic Comput. 41 (2006), no. 1, 1–29. MR2194883 (2006h:65005)

[134] Philippe Flajolet, Wojciech Szpankowski, and Brigitte Vallée, Hidden word statistics, J. ACM 53 (2006), no. 1, 147–183. MR2212002 (2007j:68190)

[135] Philippe Flajolet, Markus Nebel, and Helmut Prodinger, The scientific works of Rainer Kemp (1949–2004), Theoret. Comput. Sci. 355 (2006), no. 3, 371–381. MR2211614

[136] Philippe Flajolet, The ubiquitous digital tree, STACS 2006, Lecture Notes in Comput. Sci., vol. 3884, Springer, Berlin, 2006, pp. 1–22. MR2249355 (2007c:68021)

[137] Philippe Flajolet, Eric Fusy, Xavier Gourdon, Daniel Panario, and Nicolas Pouyanne, A hybrid of Darboux’s method and singularity analysis in combinatorial asymptotics, Electron. J. Combin. 13 (2006), no. 1, Research Paper 103, 35. MR2274318 (2008d:33003)

[138] Philippe Flajolet, Philippe Dumas, and Vincent Puyhaubert, Some exactly solvable models of urn process theory, Fourth Colloquium on Mathematics and Computer Science Algorithms, Trees, Combinatorics and Probabilities, Discrete Math. Theor. Comput. Sci. Proc., AG, Assoc. Discrete Math. Theor. Comput. Sci., Nancy, 2006, pp. 59–118. MR2509623 (2011b:60030)

[139] Philippe Flajolet, Analytic combinatorics—a calculus of discrete structures, Proceedings of the Eighteenth Annual ACM-SIAM Symposium on Discrete Algorithms, ACM, New York, 2007, pp. 137–148. MR2367438

[140] Philippe Flajolet, Éric Fusy, Olivier Gandouet, and Frédéric Meunier, HyperLogLog: the analysis of a near-optimal cardinality estimation algorithm, 2007 Conference on Analysis of Algorithms, AofA 07, Discrete Math. Theor. Comput. Sci. Proc., AH, Assoc. Discrete Math. Theor. Comput. Sci., Nancy, 2007, pp. 127–145. MR2509517 [141] Philippe Flajolet, Éric Fusy, and Carine Pivoteau, Boltzmann sampling of unlabelled struc-

tures, Proceedings of the Ninth Workshop on Algorithm Engineering and Experiments and the Fourth Workshop on Analytic Algorithmics and Combinatorics, SIAM, Philadelphia, PA, 2007, pp. 201–211. MR2498128

[142] Philippe Flajolet and Linas Vepstas, On differences of zeta values, J. Comput. Appl. Math. 220 (2008), no. 1-2, 58–73. MR2444154 (2009i:11102)

[143] Nicolas Broutin and Philippe Flajolet, The height of random binary unlabelled trees, Fifth Colloquium on Mathematics and Computer Science, Discrete Math. Theor. Comput. Sci. Proc., AI, Assoc. Discrete Math. Theor. Comput. Sci., Nancy, 2008, pp. 121–134. MR2508782

[144] Philippe Flajolet and Thierry Huillet, Analytic combinatorics of the Mabinogion urn, Fifth Colloquium on Mathematics and Computer Science, Discrete Math. Theor. Comput. Sci. Proc., AI, Assoc. Discrete Math. Theor. Comput. Sci., Nancy, 2008, pp. 549–571. MR2508814 (2010i:60032)

[145] Philippe Flajolet and Robert Sedgewick, Analytic combinatorics, Cambridge University Press, Cambridge, 2009. MR2483235 (2010h:05005)

[146] Miklós Bóna and Philippe Flajolet, Isomorphism and symmetries in random phylogenetic trees, J. Appl. Probab. 46 (2009), no. 4, 1005–1019. MR2582703 (2011b:60029)

[147] Philippe Flajolet, Stefan Gerhold, and Bruno Salvy, Lindelöf representations and (non- )holonomic sequences, Electron. J. Combin. 17 (2010), no. 1, Research Paper 3, 28. MR2578896 (2011d:11023)

[148] Roland Bacher and Philippe Flajolet, Pseudo-factorials, elliptic functions, and continued fractions, Ramanujan J. 21 (2010), no. 1, 71–97. MR2593256 (2011a:33051)

[149] Nicholas R. Beaton, Philippe Flajolet, and Anthony J. Guttmann, The unusual asymptotics of three-sided prudent polygons, J. Phys. A 43 (2010), no. 34, 342001, 10. MR2670511 (2011h:82024)